OpenStudy (precal):

related rates, calculus problem, use a^2+b^c=c^2 c = 17 a^2+b^2=289 take the derivative with respect to t

OpenStudy (anonymous):

Can you follow precal's instruction?

OpenStudy (anonymous):

I've tried this 3 times and I've gotten 4.8 but its not correct

OpenStudy (anonymous):

Show me how you take derivative:

OpenStudy (anonymous):

i've done a=9 b= square root 208 c = 17 da/dt=? db/dt=-3 dc/dt=0 a(da/dt)+b(db/dt)=c(dc/dt) 9(da/dt)+ square root 208 * (-3) = 17 * 0 9(da/dt) = 43.2666 da/dt = 4.8

OpenStudy (anonymous):

How about try: - 4.8

OpenStudy (anonymous):

that isn't right either

OpenStudy (anonymous):

Sometimes just some round off!

OpenStudy (anonymous):

There's nothing wrong with your logic and calculation !

OpenStudy (anonymous):

Did you try 4.81 (or 4.807)?

OpenStudy (anonymous):

@precal What's your result?

OpenStudy (precal):

oh I did not even do the problem at all. But give me a minute

OpenStudy (precal):